40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 266 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  3692   Mon Oct 11 22:04:28 2010 yutaUpdateCDSdamping for MCs are basically working

Background:
 Even if we can't use the Dataviewer to get the time series of each 4 DOF displacements, we can still use StripTool to monitor the ringdowns and see if the damping servo is working.

What I did:
1. Excited vibration by kicking the mirror randomly (by putting some offsets randomly and turing the filters on and off randomly).

2. Turned all the servo off by clicking "ShutDown" button.

3. Turned all the servo on by clicking "Normal" button.

3. Monitored each 4 DOF displacements with StripTool and see if there are any considerably low-Q ringdown after turning on the servo.
 The values I monitored are as follows;
  C1:SUS_MCX_SUSPOS_INMON
  C1:SUS_MCX_SUSPIT_INMON
  C1:SUS_MCX_SUSYAW_INMON
  C1:SUS_MCX_SDSEN_INMON  (X=1,2,3)

All the settings I used for this observation are automatically saved here;
  /cvs/cds/caltech/burt/autoburt/snapshots/2010/Oct/11/21:07/c1mcs.epics

Result:
 Attached is the screenshots of StripTool Graph window for each 3 MCs.
 As you can see, the dampings for each DOF, each MCs are basically working.

Note:
 Do NOT turn off all the damping servo by clicking "Damp" buttons or setting the SUSXXX_GAIN to 0. It crashes c1mcs.

Next work:
 - check and relate the signal sign with the actual moving direction of the optics
 - fix data aquisition system
 - measure Q-values when the servo is on and off using DAQ and Dataviewer

Attachment 1: SUS-MC1.png
SUS-MC1.png
Attachment 2: SUS-MC2.png
SUS-MC2.png
Attachment 3: SUS-MC3.png
SUS-MC3.png
  3691   Mon Oct 11 20:52:00 2010 ranaUpdateCDSActivation of DAQ channels for 8 optics

Bah!  We tried to get some data but totally failed. It seems like there is some rudimentary functionality in the FE process of the MC, but no useful DAQ at all.

Neither Dataview or DTT had anything. We looked and the NDS process and the DAQD process were not running on FB.

I restarted them both (./daqd -c daqdrc) & (./nds pipe > nds.log) but couldn't get anything.

There's a bunch of errors in the daqd.log like this:

CA.Client.Exception...............................................
    Warning: "Identical process variable names on multiple servers"
    Context: "Channel: "C1:SUS-MC1_SUSPOS_INMON", Connecting to: c1susdaq:57416, Ignored: c1sus.martian:57416"
    Source File: ../cac.cpp line 1208
    Current Time: Mon Oct 11 2010 18:25:15.475629328
..................................................................
CA.Client.Exception...............................................
    Warning: "Identical process variable names on multiple servers"
    Context: "Channel: "C1:SUS-MC1_SUSPIT_INMON", Connecting to: c1susdaq:57416, Ignored: c1sus.martian:57416"
    Source File: ../cac.cpp line 1208
    Current Time: Mon Oct 11 2010 18:25:15.475900124

  3690   Mon Oct 11 17:31:44 2010 yutaUpdateCDSActivation of DAQ channels for 8 optics

(Joe, Yuta)

Background:
 We need DAQ channels activated to measure Q-values of the ringdowns for each DOF, each optics with the Dataviewer.

What we did:
1. Activated the following DAQ using daqconfig (in /cvs/cds/rtcds/caltech/c1/scripts).
 C1:SUS-XX_AASEN_IN1
 C1:SUS-XX_SUSBBB_IN1
 C1:RMS-YYY_AASEN_IN1
 C1:RMS-YYY_SUSBBB_IN1
 C1:MCS-ZZZ_AASEN_IN1
 C1:MCS-ZZZ_SUSBBB_IN1
  (XX=BS,ITMX,ITMY  YYY=PRM,SRM  ZZZ=MC1,MC2,MC3  AA=UL,UR,LR,LL,SD  BBB=POS,PIT,YAW)

2. Set datarate to 2048 for each DAQ.

3. Restarted fb(frame builder).

Result:
 We succeeded in making DAQ channels appear in the Dataviewer signal list, but we can't start the measurement because c1mcs is still flaky.

Note:
 We found that c1mcs crashes everytime when turning off all the damping servo (using "Damp" buttons on the medm screen).
 It doesn't crash when all the filters are off.

  3689   Mon Oct 11 16:09:10 2010 yutaSummarySUScurrent OSEM outputs

Background:
 The output range of the OSEM is 0-2V.
 So, the OSEM output should fluctuate around 1V.
 If not, we have to modify the position of it.

What I did:
 Measured current outputs of the 5 OSEMs for each 8 suspensions by reading sensor outputs(C1:SUS-XXX_YYPDMON) on medm screens.

Result:

  BS ITMX ITMY PRM SRM MC1 MC2 MC3
UL 1.20 0.62 1.69 1.18 1.74 1.25 0.88 1.07
UR 1.21 0.54 1.50 0.99 1.77 1.64 1.46 0.31
LR 1.39 0.62 0.05 0.64 2.06 1.40 0.31 0.19
LL 1.19 0.88 0.01 0.64 1.64 1.00 0.05 1.03
SD 1.19 0.99 0.97 0.79 1.75 0.71 0.77 0.93

 White: OK (0.8~1.2)
 Yellow: needs to be fixed
 Red: BAD. definitely need fix

  3688   Mon Oct 11 10:51:36 2010 steveUpdateSUSOSEMs, OSEMs, OSEMs...those lovely little OSEMs
Attachment 1: 40dll.jpg
40dll.jpg
  3687   Mon Oct 11 10:49:03 2010 josephbUpdateCDSc1sus stability

Taking a look at the c1sus machine, it looks as if all of the front end codes its running (c1sus - running BS, ITMX, ITMY, c1mcs - running MC1, MC2, MC3, and c1rms - running PRM and SRM) worked over the weekend.  As I see no

Running dmesg on c1sus reports on a single long cycle on c1x02, where it took 17 microseconds (~15 microseconds i maximum because the c1x02 IOP process is running at 64kHz).

Both the c1sus and c1mcs models are running at around 39-42 microseconds USR time and 44-50 microseconds CPU time.  It would run into problems at 60-62 microseconds.

Looking at the filters that are turned on, it looks as it these models were running with only a single optic's worth of filters turned on via the medm screens.  I.e. the MC2 and ITMY filters were properly set, but not the others.

The c1rms model is running at around 10 microseconds USR time and 14-18 microseconds CPU time.  However it apparently had no filters on.

It looks as if no test points were used this weekend.  We'll turn on the rest of the filters and see if we start seeing crashes of the front end again.

Edit:

The filters for all the suspensions have been turned on, and all matrix elements entered.  The USR and CPU times have not appreciably changed.  No long cycles have been reported through dmesg on c1sus at this time.  I'm going to let it run and see if it runs into problems.

  3686   Sun Oct 10 18:28:25 2010 kiwamuSummarySUSITMX OSEM offsets

Because of the in-vac work on Oct. 4th (see this entry) , ITMX's OSEM offsets were changed.

The two upper OSEMs are still fine, but LL and LR seem to be out of the OSEM's range. 

The plot below shows the trends of LL's and LR's readouts for about two weeks. (The channel name are in the old convention, i.e. ITMY)

OSEM.png

 Some data were missing due to the upgrade of the frame builder.

 It is apparent that the offsets are changed after the in-vac work on Oct. 4th, and now they just show almost zero numbers.

 

The damping of ITMX can still work, if LL and LR are disabled.

At some point before pumping down, we have to check the leveling of the ITMX table again.

  3685   Sun Oct 10 18:09:02 2010 KojiSummaryCOCPhase map interferometer calibration for the data on Oct 8th, 2010

Summary:

Calibration of the phase map interferometer was calculated for the data on Oct 8th, 2010.
The calibration value is 0.1905 mm/pixel.

This is slightly smaller than the assumed value in the machine that is 0.192mm/pixel.
This means that the measured radii of curvature must be scaled down with this ratio.
(i.e. RoC(new) = RoC(old) / 0.1922 * 0.19052)


Motivation:

Our tagets of the phasemap measurement are:

1. Measure the figure errors of the mirrors
2. Measure the curvature of the mirrors

The depth of the mirror figure is calibrated by the wavelength of the laser (1064nm).
However, the lateral scale of the image is not calibrated.
Although Zygo provides the initial calibration as 0.192 mm/pixel, we should measure the calibration by ourselves.

Method:

We found an aperture mask with a grid of holes with 2mm diameter and 3mm spacing (center-to-center).
Take the picture of this aperture and calibrate the pixel size. The aperture is made of stainless and makes not interference
with the reference beam. Thus we put a dummy mirror behind the aperture. (UPPER LEFT plot)

As the holes are aligned as a grid, the FFT of the aperture image shows peaks at the corresponding pitches. (UPPER MIDDLE plot)
As the aperture was slightly rotated, the grids of the peaks are also slanted.

We can obtain the spacing of the peak grids. How can we can that values precisely? I decided to make an artificial mask image.
The artificial mask (LOWER LEFT plot) has the similar FFT pattern (LOWER MIDDLE plot). The inner product of the two
FFT results (i.e. Sum[abs(fft1) x abs(fft2)]), quite a large value is obtained if the grid pitch and the aperture angle agrees between those images.
Note that the phase information was discarded. The estimated grid spacing of the artificial mask can be mathematically obtained.

Result:

The grid pitch and the angle were manually set as initial values. Then the parameters to give the local maximum was obtained by fminsearch of Matlab.
UPPER RIGHT and LOWER RIGHT plots show the scan of the evaluation function by changing the angle and the pitch. They behave quite normal.

The obtained values are

Grid pitch: 15.74 pixel
Angle: 1.935 deg

As the grid pitch is 3mm, the calibration is

3 mm / 15.74 pixel = 0.1905 mm/pixel

Discussion:

A spherical surface can be expressed as the following formula:

z = R - R Sqrt(1-r2/R2)      (note: this can be expanded as r2/(2 R)+O(r3) )

Here R is the RoC and r is the distance from the center. This means that the calibration of r quadratically changes the curvature.
We have measured the RoC of the spare SRM. We can compare the RoCs measured by this new metrology IFO and the old one,
as well as the one by Coastline optics. 

 

Attachment 1: calibration.pdf
calibration.pdf
  3684   Sun Oct 10 16:59:20 2010 KojiOmnistructureCOCPRM phase map measurement at Downs SB 014

[Kiwamu, Yuta, Koji]

We went to the new metrology lab at Downs subbasement (Rm014) in order to measure the phase map of the delivered PRMs.

It's brand-new. So we had to measure the reference phase map, calibration as well as the phase map of our mirrors (3 PRMs and 1 spare SRM). It took a whole day...

Attachment 1: IMG_3646.jpg
IMG_3646.jpg
Attachment 2: IMG_3647.jpg
IMG_3647.jpg
  3683   Sun Oct 10 16:44:59 2010 KojiOmnistructurePhotosKepco Tube HV supply
Attachment 1: IMG_3637.jpg
IMG_3637.jpg
Attachment 2: IMG_3640.jpg
IMG_3640.jpg
  3682   Fri Oct 8 17:36:16 2010 steveFrogsPhotosvisiting undergrads

Prof Alan Weistein guided the 24 student through the 40m. His performance was rated as  an enthusiastic 9.5

Attachment 1: P1060916.JPG
P1060916.JPG
Attachment 2: P1060921.JPG
P1060921.JPG
Attachment 3: P1060922.JPG
P1060922.JPG
Attachment 4: P1060915.JPG
P1060915.JPG
  3681   Fri Oct 8 17:35:24 2010 josephbUpdateCDSstatus of c1ioo, c1sus and rfm

RFM is still not working.  I can see data on a filter just before it reaches the RFM sending code, but I see only zeros on the receiving side.

c1sus machine and c1x02, c1sus, c1mcs, c1rms are running.  At the moment, the c1mcs model is running at about 42 microseconds for USR time and 56 microseconds for CPU MAX, which is close to the 61 microsecond limit.  This is with MC filters on.  So far it has not been late, but its not clear to me if its going to stay that way.  So far I haven't been able to isolate why it sometimes slows down after a few minutes.  Also, it was running faster earlier in the day (around 30-ish microseconds) and I believe it has slowed down slightly in the last hour or two.

c1ioo machine and c1x03, c1ioo are running. However its not doing very much good as I can't get any data transferred from it to any of the optic suspensions. I need to spend some more time debugging this and then grab Alex I think.

  3680   Fri Oct 8 15:57:32 2010 josephbUpdateCDSc1ioo status

I've been trying to figure out why the c1ioo machine crashes when I try to run the c1ioo front end.

I tried removing some RFM components from the c1ioo model, and then the c1gpt model (Kiwamu's green locking model) as an easier test case.  Both cause the machine to lock up once they start running.  Lastly, I tried running the c1x02 and c1sus models on the c1ioo machine instead of the c1sus machine, after first turning off all models on c1sus.  This led to the same lockup. 

Since those models run fine on the c1sus machine, I could only conclude that a recent change in the fe code generator or the Gentoo kernel and the Sun X4600 computer don't work together at the moment.

After talking to Alex, he got the idea to check if the monitor() and mwait() were supported on the c1ioo machine.  These function calls were added relatively recently, and are   used to poll a memory location to see if something has been written there, and then do something when it is.  Apparently the Sun X4600 computers do not support this call.  Alex has modified the code to not use these functions calls, at least for now.  He'd like to add a check to the code so it does use those calls on machines which have them supported.

After this change, the c1ioo and c1gpt front end codes do in fact run.

  3679   Fri Oct 8 12:29:21 2010 KevinUpdateComputersNew Netgear Switch

I removed some old equipment from the rack outside the control room and stacked them next to the filing cabinets in the control room. I also mounted the new Netgear switch in the rack.

  3678   Fri Oct 8 12:21:11 2010 josephbUpdateCDSchecking MC1 suspension damping

Upon investigation, it appears that the c1mcs model was (and still is) timing out after a random amount of time. Or in other words, it at some point it was taking too long to do all the calculations for a single cycle and falling behind. The evidence for this is from the dmesg command when run on c1sus.

There's a bunch of lines like:

[ 8877.438002] c1mcs: cycle 568 time 62; adcWait 0; write1 0; write2 0; longest write2 0

[ 8877.438002] c1mcs: cycle 569 time 62; adcWait 0; write1 0; write2 0; longest write2 0

With a final line like: [ 8877.439003] c1mcs: ADC TIMEOUT 1 2405 37 2277

This last line indicates in fell so far behind it gave up.

However, this doesn't actually seem to be related to the amount of computation going on with the front end. I restarted the c1mcs model this morning by logging into the c1sus machine, and changing to the /opt/rtcds/caltech/c1/target/c1mcs/bin directory and running:

lsmod

sudo rmmod c1mcsfe

sudo insmod c1mcsfe.ko

The first line just lists the running modules. The second removes the c1mcs module, and the third starts it up again.

I proceeded to turn all the filters and and set all the matrix values while keeping an eye on the C1MCS-GDS_TP.adl screen and its timing indicator. It was running fine with all these turned on. Then I ran a dtt session from rosalba (going to /opt/apps/, typing bash, then source gds-env.bash, and finally diaggui) as well as a dataviewer and looked at 6 test point channels. It received data fine.

However, about 2 minutes after I had stopped doing this (although the dataviewer realtime session was still going) the USR timing jumped from about 20 microseconds to 35 microseconds, and the CPU Max timing jumped to the 50-60 microsecond range. At this point dmesg started reporting things like:

[54143.465613] c1mcs: cycle 1076 time 62; adcWait 0; write1 0; write2 0; longest write2 0

[54143.526004] c1mcs: cycle 2064 time 62; adcWait 0; write1 0; write2 0; longest write2 0

About a minute later the code gave up and reported a ADC timeout via dmesg. None of the other front ends seem to be affected.

I had to clear the test points manually after stopping dataviewer and dtt by going to rosalaba,using the sourced gds-env.bash, and running diag -l. I then typed "tp clear 36 *" to clear all the test points on the model with FEC number 36 (corresponding to c1mcs).

I removed and restarted c1mcs again, trying to turn on a few things at a time and letting it run for a few minutes to see if I could narrow down if its one particular filter perhaps reaching an underflow and starting to bog down the computations. However, after about 45 minutes of this, the model is still running and I've turned all the filters on and have been running about 8 test points with no problem, so the problem is clearly intermittent.

Quote:

Notes:
 c1mcs crashed many times during the investigation, and I had to kill and restart it again and again.
 It seemed to be easily crashed when filters are on, and so I couldn't check whether the damping servo is working correctly or not today.

Next work:

  - fix c1mcs (and maybe others)
  - check the damping servo by comparing the displacements of each 4 degrees of freedom when servo in off and on.

 

  3677   Fri Oct 8 10:38:03 2010 steveConfigurationSAFETY2W laser shutter is closed

Quote:

The 2 W Innolight shutter is closed and enclosure door removed. Beam path blocked. Do not change this condition.

The PSL output beam guide was upgraded from 2" to 8" OD . It is green ready. Shutter is open.

Attachment 1: P1060905.JPG
P1060905.JPG
Attachment 2: P1060910.JPG
P1060910.JPG
  3676   Fri Oct 8 07:41:42 2010 steveConfigurationSAFETY2W laser shutter is closed

The 2 W Innolight shutter is closed and enclosure door removed. Beam path blocked. Do not change this condition.

  3675   Thu Oct 7 23:24:44 2010 yutaUpdateCDSchecking MC1 suspension damping

Background:
 The new CDS is currently being set up.
 We want to see if the damping servo of the suspensions are working correctly.
 But before that, we have to see if the sensors and the coils are working correctly.
 Among the 8 optics, MCs take top priority because of the green beam. for the alignment of the in-vac optics.

What I did:
 By seeing the 5 sensor outputs (C1:SUS-MC1_XXSEN_IN1, XX=UL,UR,LR,LL,SD) with the Data Viewer, I checked if all the coils are kicking in the supposed direction and all the sensors are sensing that kick correctly.

 All the matrices elements were set to the ideal values(-1 or 0 or 1) this time.

Result:
 They were perfect.
1. POSITION seemed to be POSITION
 When the offset(C1:SUS-MC1_SUSPOS_OFFSET) was added, all the sensor output moved to the same direction.
2. PITCH seemed to be PITCH
 When the offset(C1:SUS-MC1_PIT_COMM) was changed, UL&UR and LL&LR went to the different direction.
3. YAW seemed to be YAW
 When the offset(C1:SUS-MC1_YAW_COMM) was changed, UL&LL and UR&LR went to the different direction.
4. SIDE seemed to be SIDE
 When the offset(C1:SUS-MC1_SDSEN_OFFSET) was added, DC level of the SD sensor output changed.

Notes:

 c1mcs crashed many times during the investigation, and I had to kill and restart it again and again.
 It seemed to be easily crashed when filters are on, and so I couldn't check whether the damping servo is working correctly or not today.

Next work:

  - fix c1mcs (and maybe others)
  - check the damping servo by comparing the displacements of each 4 degrees of freedom when servo in off and on.

  3674   Thu Oct 7 17:53:07 2010 yutaUpdateComputersFarfalla with Firefox, fixed

(Kiwamu, Yuta)

Symptom:
 Farfalla(Acer Aspire one KAV60 netbook) couldn't run Firefox and returned the following error:
  Error: platform version '1.9.2.8' is not compatible with
  minVersion >= 1.9.2.9
  maxVersion <= 1.9.2.9

What we did:
1. Added the following line to ~/.cshrc.
  alias firefox "/usr/lib/firefox-3.6/firefox"
2. Made a cool launcher for firefox.

Result:
 Farfalla can fly into the web with Firefox now.

Notes:

 Even if it was then before, bash could run firefox. tsch couldn't.
 The command firefox was /cvs/cds/caltech/apps/linux/bin/firefox for tsch, because of the source /cvs/cds/caltech/cshrc.40m.
 I did this to zita which had the same symptom, too.

Next work:

 Farfalla wakes up on the wrong side of the bed. This has to be fixed.

  3673   Thu Oct 7 17:19:55 2010 josephb, alex, rolfUpdateCDSc1sus status

As noted by Koji, Alex and Rolf stopped by.

We discussed the feasibility of getting multiple models using the same DAC.  We decided that we infact did need it. (I.e. 8 optics through 3 DACs does not divide nicely), and went about changing the controller.c file so as to gracefully handle that case.  Basically it now writes a 0 to the channel rather than repeating the last output if a particular model goes down that is sharing a DAC.

In a separate issue, we found that when skipping DACs  in a model (say using DACs 1 and 2 only) there was a miscommunication to the IOP, resulting in the wrong DACs getting the data.  the temporary solution is to have all DACs in each model, even if they are not used.  This will eventually be fixed in code.

At this point, we *seem* to be able to control and damp optics.  Look for a elog from Yuta confirming or denying this later tonight (or maybe tomorrow).

 

  3672   Thu Oct 7 16:34:06 2010 yutaUpdatePSLmeasured PMC's laser power-output relation

Result2:
 Attached is the visibility vs incident power(assuming output of the PD is proportional to the input laser power).
 Ideally, the graph should be flat. (In another words, attached graph in the elog #3667 shoud be linear.)
 But the visibility reduces with higher laser power in this graph. This is maybe because of the thermal effect. I'm thinking about how to confirm this.

 Quote:

It was a bit difficult to comprehend the result.
Is it good? or bad? Have you seen the thermal effect? or not?

- Put linear lines to show the visibility of the cavity.

- Calibrate the incident power and make another plot to show the visibility (%) vs the incident power (W).

Quote:

(Rana, Yuta)

Motivation:

 We wanted to see thermal effects on the PMC.

What I did yesterday:
 Changed the current of the NPRO from 2A to 0.8A and measured the power of the reflected/transmitted light from the PMC when locked.
 I also measured the power of the reflected light when PMC is not locked (It supposed to be proportional to the output power of the laser).

Result:
 Attached. Hmmmm......
 At several points of the laser current, I could'nt lock the PMC very well. The power of the reflected/transmitted light depend on the offset voltage of the PZT.
 When the laser power was weak(~<0.9A), the power of reflected/transmitted light changed periodically(~ several minutes).

 

 

Attachment 1: PMCvis.png
PMCvis.png
  3671   Thu Oct 7 16:21:02 2010 KojiOmnistructureCDSBig 3

Both Rolf and Alex (at least his elbow) together visited the 40m to talk with Joe for the CDS.

40m is the true front line of the CDS development!!!

Attachment 1: IMG_3642.jpg
IMG_3642.jpg
  3670   Thu Oct 7 15:21:51 2010 steveUpdatePSLRF filter for PMC

We got two small RF filter for the PMC from Valera  They are made by http://www.larkengineering.com/   "MC35.5-3-AB" sma, 29300-01

Attachment 1: 35.5MHZ.pdf
35.5MHZ.pdf
  3669   Thu Oct 7 15:05:46 2010 KojiUpdatePSLmeasured PMC's laser power-output relation

It was a bit difficult to comprehend the result.
Is it good? or bad? Have you seen the thermal effect? or not?

- Put linear lines to show the visibility of the cavity.

- Calibrate the incident power and make another plot to show the visibility (%) vs the incident power (W).

Quote:

(Rana, Yuta)

Motivation:

 We wanted to see thermal effects on the PMC.

What I did yesterday:
 Changed the current of the NPRO from 2A to 0.8A and measured the power of the reflected/transmitted light from the PMC when locked.
 I also measured the power of the reflected light when PMC is not locked (It supposed to be proportional to the output power of the laser).

Result:
 Attached. Hmmmm......
 At several points of the laser current, I could'nt lock the PMC very well. The power of the reflected/transmitted light depend on the offset voltage of the PZT.
 When the laser power was weak(~<0.9A), the power of reflected/transmitted light changed periodically(~ several minutes).

 

  3668   Thu Oct 7 14:57:52 2010 josephb, yutaUpdateCDSc1sus status

Around noon, Yuta and I were trying to figure out why we were getting no signal out to the mode cleaner coils.  It turns out the mode cleaner optic control model was not talking to the IOP model. 

Alex and I were working under the incorrect assumption that you could use the same DAC piece in multiple models, and simply use a subset of the channels.  He finally went and asked Rolf, who said that the same DAC simulink piece in different models doesn't work.  You need to use shared memory locations to move the data to the model with the DAC card.  Rolf says there was a discussion (probably a long while back) where it was asked if we needed to support DAC cards in multiple models and the decision was that it was not needed.

Rolf and Alex have said they'd come over and discuss the issue.

In the meantime, I'm moving forward by adding shared memory locations for all the mode cleaner optics to talk to the DAC in the c1sus model.

 

Note by KA: Important fact that is worth remembering

  3667   Thu Oct 7 14:39:50 2010 yutaUpdatePSLmeasured PMC's laser power-output relation

(Rana, Yuta)

Motivation:

 We wanted to see thermal effects on the PMC.

What I did yesterday:
 Changed the current of the NPRO from 2A to 0.8A and measured the power of the reflected/transmitted light from the PMC when locked.
 I also measured the power of the reflected light when PMC is not locked (It supposed to be proportional to the output power of the laser).

Result:
 Attached. Hmmmm......
 At several points of the laser current, I could'nt lock the PMC very well. The power of the reflected/transmitted light depend on the offset voltage of the PZT.
 When the laser power was weak(~<0.9A), the power of reflected/transmitted light changed periodically(~ several minutes).

Attachment 1: PMCreflect.png
PMCreflect.png
  3666   Thu Oct 7 10:48:41 2010 josephb, yutaUpdateCDSc1sus status

This problem has been resolved.

Apparently during one of Alex's debugging sessions, he had commented out the feCode function call on line 1532 of the controller.c file (located in /opt/rtcds/caltech/c1/core/advLigoRTS/src/fe/ directory).

This function is the one that actually calls all the front end specific code and without it, the code just doesn't do any computations.  We had to then rebuild the front end codes with this corrected file.

Quote:

At the moment, c1sus and c1mcs on the c1sus machine seem to be dead in the water.  At this point, it is unclear to me why.

Apparently during the 40m meeting, Alex was able to get test points working for the c1mcs model.  He said he "had to slow down mx_stream startup on c1sus".   When we returned at 2pm, things were running fine. 

We began updating all the matrix values on the medm screens.  Somewhere towards the end of this the c1sus model seemed to have crashed, leaving only c1x02 and c1mcs running.  There were no obvious error messages I saw in dmesg and the target/c1sus/logs/log.txt file (although that seems to empty these days).  We quickly saved to burt snap shots, one of c1sus and one of c1mcs and saved them to /opt/rtcds/catlech/c1/target/snapshots directory temporarily.  We then ran the killc1sus script on c1sus, and then after confirming the code was removed, ran the startup script, startc1sus.  The code seemed to come back partly.  It was syncing up and finding the ADC/DAC boards, but not doing any real computations.  The cycle time was reporting reasonably, but the usr time (representing computation done for the model) was 0.  There were no updating monitor channels on the medm screens and filters would not turn on.

At this point I tried bringing down all 3 models, and restarting c1x02, then c1sus and c1mcs.  At this point, both c1sus and c1mcs came back partly, doing no real calculations.  c1x02 appears to be working normally (or at least the two filter banks in that model are showing changing channels from ADCs properly).  I then tried rebooting the c1sus machine.  It came back in the same state, working c1x02, non-calculating c1sus and c1mcs.

 

  3665   Thu Oct 7 10:37:42 2010 josephbUpdateCDSc1sus with flaky ssh

Currently trying to understand why the ssh connections to c1sus  are flaky.  This morning, every time I tried to make the c1sus model on the c1sus machine, the ssh session would be terminated at a random spot midway through the build process.  Eventually restarting c1sus fixed the problem for the moment.

However, previously in the last 48 hours, the c1sus machine had stopped responding to ssh logins while still appearing to be running the front end code.  The next time this occurs, we should attach a monitor and keyboard and see what kind of state the computer is in.  Its interesting to note we didn't have these problems before we switched over to the Gentoo kernel from the real-time linux Centos 5.5 kernel.

  3664   Thu Oct 7 08:39:39 2010 steveBureaucracySAFETYJoonho receives safety training

Our new undergrad Joonho Lee received 40m specific basic safety training yesterday.

Yuta and Joonho are scheduled to participate in the LIGO-laser safety tutelage with Peter K on Oct 12

  3663   Wed Oct 6 22:46:36 2010 kiwamuUpdateGreen LockingSHG at PSL table

 I put some optics to get the green SHG on the PSL table.

Now a green light successfully comes out from the doubling crystal.

Since the optical layout of the PSL table was dramatically changed, I had to re-setup the green SHG stuff. 

 

 - what I did

I put a steering mirror after the 90/10% pick off mirror, and then a half wave plate and a convex lens.

The lens is approximately on the right place.

 To get the green beam easily, temporarily I raised the current of the NPRO up to 2 A.

I connected the oven to the heat controller, set the temperature to 36.8 deg which is the set point previously used.

After putting and aligning the oven, I finally obtained the green beam.

At the end of the work I set the NPRO current back to 0.9 A and relocked the PMC.

 

- things to be done

 1. more precise mode matching

 2. optimization of the temperature 

  3662   Wed Oct 6 16:16:48 2010 josephb, yutaUpdateCDSc1sus status

At the moment, c1sus and c1mcs on the c1sus machine seem to be dead in the water.  At this point, it is unclear to me why.

Apparently during the 40m meeting, Alex was able to get test points working for the c1mcs model.  He said he "had to slow down mx_stream startup on c1sus".   When we returned at 2pm, things were running fine. 

We began updating all the matrix values on the medm screens.  Somewhere towards the end of this the c1sus model seemed to have crashed, leaving only c1x02 and c1mcs running.  There were no obvious error messages I saw in dmesg and the target/c1sus/logs/log.txt file (although that seems to empty these days).  We quickly saved to burt snap shots, one of c1sus and one of c1mcs and saved them to /opt/rtcds/catlech/c1/target/snapshots directory temporarily.  We then ran the killc1sus script on c1sus, and then after confirming the code was removed, ran the startup script, startc1sus.  The code seemed to come back partly.  It was syncing up and finding the ADC/DAC boards, but not doing any real computations.  The cycle time was reporting reasonably, but the usr time (representing computation done for the model) was 0.  There were no updating monitor channels on the medm screens and filters would not turn on.

At this point I tried bringing down all 3 models, and restarting c1x02, then c1sus and c1mcs.  At this point, both c1sus and c1mcs came back partly, doing no real calculations.  c1x02 appears to be working normally (or at least the two filter banks in that model are showing changing channels from ADCs properly).  I then tried rebooting the c1sus machine.  It came back in the same state, working c1x02, non-calculating c1sus and c1mcs.

  3661   Wed Oct 6 15:56:14 2010 josephbHowToCDSHow to load matrices quickly and easily

Awhile ago I wrote several scripts for reading in medm screen matrix settings and then writing them out.  It was meant as kind of a mini-burt just for matrices for switching between a couple of different setups quickly.

Yuta has expressed interest in having this instruction available.

In /cvs/cds/caltech/scripts/matrix/ there are 4 python scripts:

saveMatrix.py, oldSaveMatrix.py, loadMatrix.py, oldLoadMatrix.py

The saveMatrix.py and loadMatrix.py are for use with the current front end codes (which start counting from 1), where as the old*.py files are for the old system.

To use saveMatrix.py you need to specify the number of inputs, outputs, and the base name of the matrix (i.e. C1:LSP-DOF2PD_MTRX is the base of C1:LSP-DOF2PD_MTRX_0_0_GAIN for example), as well as an ouptut file where the values are stored.

So to save the BS in_matrix setting you could do (from /cvs/cds/caltech/scripts/matrix/)

./saveMatrix.py -i 4 -o 5 -n "C1:SUS-BS_TO_COIL_MTRX" -f -d ./to_coil_mtrx.txt

The -i 4 indicates 4 inputs, the -o 5 indicates 5 outputs, -n "blah blah" indicates the base channel name, -f indicates a matrix bank of filters (if its just a normal matrix with no filters, drop the -f flag), and -d ./to_coil_mtrx.txt indicates the destination file.

To write the matrix, you do virtually the same thing:

./loadMatrix.py -n "C1:SUS-PRM_TO_COIL_MTRX" -f -d ./to_coil_mtrx.txt

In this case, you're writing the saved values of the BS, to the PRM.  This method might be faster if you're trying to fill in a bunch of new matrices that are identical rather than typing 1's and -1's 20 times for each matrix.

I'll have Yuta add his how-to of this to the wiki.

  3660   Wed Oct 6 14:49:54 2010 ranaSummaryloreSteve on the sea

jacques.jpg

  3659   Wed Oct 6 12:00:23 2010 josephb, yuta, kiwamuUpdateCDSFound and fixed filter sampling rate problem with suspensions

While we looking using dtt and going over the basics of its operation, we discovered that the filter sample rates for the suspensions were still set to 2048 Hz, rather than 16384 Hz which is the new front end.  This caused the filters loaded into the front ends to not behave as expected.

After correcting the sample rate, the transfer functions obtained from dtt are now looking like the bode plots from foton.

We fixed the C1SUS.txt and C1MCS.txt files in the /opt/rtcds/caltech/c1/chans/ directory, by changing the SAMPLING lines to have 16384 rather than 2048.

  3658   Wed Oct 6 11:03:51 2010 josephb, yutaHowToCDSHow to start diaggui for right now

I'm hoping to get a proper install this week done, but for now, this a stop gap.

To start diagnostic test tools, go to rosalba.  (Either sit at it, or ssh -X rosalba).

cd /opt/apps

type "bash", this starts a bash shell

source gds-env.bash

diaggui

 --------- Debugging section ------

If that throws up errors, try looking with "diag -i" and see if there's a line that starts with nds.  In the case last night, Alex had not setup a diagconf configuration file in the /etc/xinetd.d directory, which setups up the diagconf service under the xinit service.  To restart that service (if for example the nds line doesn't show up), go to /etc/init.d/ and type "sudo xinit start" (or restart).

Other problems can include awg and/or tpman not running for a particular model on the front end machine.  I.e. diag -i should show 3 results from 192.168.113.85 (c1x02, c1sus, c1mcs) at the moment , for both awg and tp.  If not, that means awg and tpman need to be restarted for those.

These can be started manually by going to the front end, to the /opt/rtcds/caltech/c1/target/gds/bin/ directory, and running awgtpman -s sysname (or in the case of IOP files [c1x02, c1x03, etc], awgtpman -s sysname -4.  Better is probably to run the start scripts which live /opt/rtcds/caltech/c1/scripts/ which kills and restarts all the process for you.

 

 

  3657   Wed Oct 6 00:32:01 2010 ranaSummaryDAQNDS2

This is the link to the NDS2 webpage:

https://www.lsc-group.phys.uwm.edu/daswg/wiki/NetworkDataServer2

We should install this so that we can use this modern interface to get 40m data from outside and inside of the 40m.

  3656   Tue Oct 5 21:22:46 2010 yutaUpdateVACgreen beam reached PSL table

YES ! We got the green light coming out from the OMC chamber to the PSL table !

 (Koji, Kiwamu, Yuta)

Background

As a result of the work in the chambers, the green beam reached the OMC chamber yesterday.
Today's mission was to let the green beam coming out from the chambers to the PSL table.

What we did:
1. Installed the last three steering mirrors.
  The mirrors were 532 HR mirror with AR and 1deg wedge (CVI Y2-LW-1-2050-UV-45-P/AR).
  Two of them are placed to the MC chamber. Another one was to the OMC chamber

During putting the mirrors to the MC chamber, we found that a cable tower was sitting on a position exactly where we wanted to put one of the mirrors.

So we moved the tower to the very south west corner. 

 

2. Installed a periscope to the MC chamber

The function of this periscope is to lower the beam height of the green light which is risen up by another periscope in the BS chamber.

We aligned it to the green beam, so that the beam hits the center of the mirrors on the periscope.  

 

3. Aligned the optics

We aligned the green mirrors so that we can let the green beam go out from the chamber.

Actually the inside of the OMC chamber didn't look like the same as that of our optical layout.

For example there is unknown base plate, which apparently disturbs the location of our last steering mirror.

Therefore we had to change the designed position of the steering mirror.

Now the mirror is sitting near the designed position (~ 1/2 inch off), but it's fine because it doesn't clip any 1064 beam.

 
Result:

The green beam is now hitting the north wall of the PSL table.

 

Notes:

The green beam looks having some fringes, it may be caused by a multiple reflection from a TT when the green beam goes through it. We are going to check it. 

 

Next work:

- Damping of the suspended optics

- Resurrect MC and its stable lock

- Remove MCT pickoff path

- Align optics in the main path

- Recycled Michelson lock

 IMG_3627.jpg

 

  3655   Tue Oct 5 18:27:18 2010 Joonho LeeSummaryElectronicsCCD cable's impedence

Today I checked the CCD cables which is connected to the VIDEOMUX.

17 cables are type of RG59, 8 cables are type of RG58. I have not figured out the type of other cables(23 cables) yet.

The reason I am checking the cables is for replacing the cables with impedance of 50 or 52 ohm by those with impedance of 75 ohm.

After I figures out which cable has not proper impedance, I will make new cables and substitute them in order to match the impedance, which would lead to better VIDEO signal.

To check the impedance of each CCD cable, I went to the VIDEOMUX and looked for the label on the cable's surface.

Type of RG59 is designated to the cable of impedance 75ohm. I wrote down each cable's input or output channel number with observation(whether it is of type RG59 or not).

The result of observation is as follows.

Type channel number where it is connected to
Type 59 in#2, in#11, in#12, in#15, in#18, in#19, in#22, in#26, out#3, out#4, out#11, out#12, out#14, out#17, out#18, out#20, out#21
Type 58 in#17, in#23, in#24, in#25, out#2, out#5, out#7, out#19
unknown type others

 

For 23 cables that I have not figured out their type, cables are too entangled so it is limited to look for the label along each cable.

I will try to figure out more tomorrow. Any suggestion would be really appreciated.

  3654   Tue Oct 5 17:39:21 2010 steveConfigurationVACwest access connector removed

Kiwamu, Koji, Yutah, Rana, Jan and Steve

We removed the west access connector this afternoon.

  3653   Tue Oct 5 16:58:41 2010 josephb, yutaUpdateCDSc1sus front end status

We moved the filters for the mode cleaner optics over from the C1SUS.txt file in /opt/rtcds/caltech/c1/chans/ to the C1MCS.txt file, and placed SUS_ on the front of all the filter names.  This has let us load he filters for the mode cleaner optics.

At the moment, we cannot seem to get testpoints for the optics (i.e. dtt is not working, even the specially installed ones on rosalba). I've asked Yuta to enter in the correct matrix elements and turn the correct filters on, then save with a burt backup.

  3652   Tue Oct 5 16:30:00 2010 josephb, yutaHowToCDSScreen settings and medm screens for new system

You can find the sitemap medm screen in

/opt/rtcds/caltech/c1/medm/master

The settings for the screens were last saved by burt in the original system on Sept 29, 2010 at 11:07.  So go to the

/cvs/cds/caltech/burt/autoburt/snapshots/2010/Sep/29/11:07

directory.  You can grep for the channels in the files in this directory.

You can also then use the autoBurt.req file in the /opt/rtcds/caltech/c1/target/sysname/sysnameepics (c1sus/c1susepics) to backup the settings entered.  Save to the /opt/rtcds/caltech/c1/target/snapshots directory for now.

 

 

  3651   Tue Oct 5 14:11:09 2010 josephb, alexUpdateCDSGoing to from rtlinux to Gentoo requires front end code clean out

Apparently when updating front end codes from rtlinux to the patched Gentoo, certain files don't get deleted when running make clean, such as the sysfe.rtl files in the advLigoRTS/src/fe/sys directories.  This fouls the start up scripts by making it think it should be configured for rtlinux rather than the Gentoo kernel module.

  3650   Tue Oct 5 13:59:17 2010 Leo SingerConfigurationComputersInstalled x264-devel on Allegra

 I installed the package x264-devel on allegra.martian.  This package provides headers and libraries for the popular h264 video codec.  I am going to use this in the GStreamer streaming media server on Allegra.

  3649   Tue Oct 5 13:52:15 2010 ranaUpdateCDSproof of trend

Untitled.png

  3648   Tue Oct 5 13:46:26 2010 josephb, alexUpdateCDSRestarted fb trending

Fb is now once again actually recording trends.

A section of the daqdrc file (located in /opt/rtcds/caltech/c1/target/fb/ directory) had been commented out by Alex and never uncommented.  This section included the commands which actually make the fb record trends.

The section now reads as:

# comment out this block to stop saving data

#
start frame-saver;
sync frame-saver;
start trender;
start trend-frame-saver;
sync trend-frame-saver;
start minute-trend-frame-saver;
sync minute-trend-frame-saver;
start raw_minute_trend_saver;
#start frame-writer "225.225.225.1" broadcast="131.215.113.0" all;
#sleep 5;

  3647   Tue Oct 5 11:42:20 2010 kiwamuSummaryGreen Lockingdeveloping green locking plant

 With a help from Joe, I made a diagram of the simulated plant for green locking in order to get better understanding and consensus.

Eventually these simulated plants will help us developing (sometimes debugging) the digital control systems.

Here is the diagram which tells us how we will setup and link the control/plant models and on which machine they will be running.

green_sm3.png

 

Basically upper side represents the realtime control, and the lower side is the simulated plant.

The models are talking to each other via either a local shared memory (orange line) or the reflective memory network (purple line).

 Each model is stil not systematically named, at some point we have to have an absolute standard for naming the models.

 

- current model names -

  GCV = Green Control model at Vertex

  GCX(Y) = Green Control model at X (Y) end

  GPV = Green Plant model at Vertex

 

 - things to be done -

1. let the RFM work

2. revise the old plant models : SUP, SPX(Y) and LSP

 

  3646   Tue Oct 5 09:26:04 2010 steveConfigurationPEMmoved accelerometers

Accelerometers xyz were moved from IOO-south/west  corner to under PSL table. They were turned off for ~20 minutes.

Guralp was also moved eastward about 2 ft. It is not leveled.

This is part of the preparation to remove access connector.

  3645   Mon Oct 4 19:48:00 2010 yutaUpdateVACseveral mirrors installed to ITMX/BS chamber

(Koji, Kiwamu, Yuta)

Lots of progress for the optics placement in the vacuum chamber.
We are ready to go to the next step: open the access connector between IMC and OMC.
 


Background:

The actual work in the vacuum chamber started.
The first goal of the vent is to get the green beam coming out from the chambers to the PSL table.


What we did:

1. Inspection of the tip-tilt suspension

The alignments of the TTs were inspected. We had five TTs having been sitting on the table in Bob's clean room.
Prior to their installation we checked the alignments of those because they sometimes showed large hysteresis mainly in the pitch directions.

If a TT has the hysteresis, the pitch position doesn't come back to the previous position. This may cause an issue of the interferometer operation.

- SN001, SN003, SN005 looked well aligned and show no hysteresis.

- The alignment of SN002 was not so good (theta ~ 0.004 rad ) compared to those three, but no hysteresis, we think this guy is still acceptable for the installation.

- SN004 had an apparent hysteresis. This guy doesn't come back to the previous place, being applied a touch. We have to fix this at some point.

2. Work in the ITMX chamber

Now all of the optic in this chamber was installed in the approximate place.

- Installed POX/POP steering mirrors.

- TT(SN003) was placed.

- The two steering optics for ITMX OPLEV were placed at the designed positions.

3. Work in the BS chamber

Installed 2 TTs to the BS chamber.
- SR3: TT(SN001)
- PR3: TT(SN005)

After the alignment of the green steering mirrors, we confirmed
the green beam is successfully hitting the west wall of the OMC chamber.

Those TTs are approximately aligned using the weakly reflected green beams.

Next work:

- Open the access connector

- Place another periscope and two steering mirrors for green

- Damping of the suspended optics

- Resurrect MC and its stable lock

- Remove MCT pickoff path

- Align optics in the main path

- Recycled Michelson lock

 

Attachment 1: P1060901.JPG
P1060901.JPG
Attachment 2: P1060904.JPG
P1060904.JPG
  3644   Mon Oct 4 15:28:10 2010 josephbUpdateCDSTrying to get c1ioo booting as Gentoo.

I modified the dhcpd.conf file in /etc/dhcp on the fb machine.  I added a entry for c1ioo, listing its MAC address and ip number near the bottom of the file.  I then restarted the dhcp server using "sudo /etc/init.d/dhcpd restart" while on the fb machine.

I also modified the rtsystab, which is used to determine which front end codes start on boot up of a machine.  I added a line: c1ioo   c1x03  c1ioo

I am now in the process of getting c1ioo to come up as a Gentoo machine so I can build a model with an RFM connection in it and test the communication between c1sus and c1ioo.  This involves removing the hard drives and checking to make sure the boot priority is correct (i.e. it checks for a network boot).

  3643   Mon Oct 4 13:48:41 2010 Leo SingerConfigurationComputersUninstalled gstreamer-devel and gstreamer-plugins-base-devel on rosalba

 I uninstalled gstreamer-devel and gst-plugins-base-devel on Rosalba.  Here is the command I ran:

$ sudo yum remove gstreamer-devel gstreamer-plugins-base-devel

 

Actually, I had installed these myself a few days earlier, before I knew that I should be recording such changes in the elog.  I'm sorry!

ELOG V3.1.3-